OpenStudy (qqstory):

There are 5 blue and 3 orange marbles in an urn. You will be picking without replacement. What is the probability of picking a blue marble on the second pick if you don't know the color of the first pick?

OpenStudy (qqstory):

answer and explain please..

OpenStudy (qqstory):

why would it be a 5/8? when its on the second draw..

OpenStudy (qqstory):

Its the part where " if you don't know the color of the first pick" is confusing me..

OpenStudy (anonymous):

yeah you are right that is confusing me as well

OpenStudy (qqstory):

lol

OpenStudy (anonymous):

it cant be 5/8 for sure it can be 5/7

OpenStudy (anonymous):

because we already picked one from the first draw right OR it can be 4/7

OpenStudy (qqstory):

so its 4/7 +5/7?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its either 5/7 or 4/7 but i dont know which one for sure actually it doesnt really specify which colour was taken out in the first try

OpenStudy (anonymous):

it can be 5/7 if orange was taken out on the first try or it can be 4/7 if blue was taken out on the first try

OpenStudy (qqstory):

o.o so i need a numberic answer.. o.o

OpenStudy (anonymous):

@hamza_b23 I agree with on 4/7 or 5/7. But I would guess that it's 4/7 because when you first picked it most likely would have been a blue.

OpenStudy (qqstory):

- - miss_deva took u 5mins to type that LOl.

OpenStudy (anonymous):

I was away for about 2 mins okay?! Lol So it technically toke me 3. - -

OpenStudy (qqstory):

lol

OpenStudy (anonymous):

:P

OpenStudy (qqstory):

Mr.@hamza_b23 what about this? What is the probability of picking a blue marble on the first pick, given that the 2nd pick is blue?

OpenStudy (anonymous):

lol but yeah we cant write just 4/7 because the other option is still awailable you know theres hardly any information

OpenStudy (anonymous):

by intuition 5/8

OpenStudy (anonymous):

\[ \frac{3\ 5}{8\ 7}+\frac{5\ 4}{8\ 7} =\frac 5 8 \]

OpenStudy (anonymous):

it is not making sense if i do it by the conditional probability

OpenStudy (anonymous):

Can you explain in words my solution in my post?

OpenStudy (anonymous):

because look at it this way if the second one is blue and the first one is blue so to pick the first blue one is 5/8 obviously and then the next one is 4/7 you see?

OpenStudy (anonymous):

sorry @eliassaab but how did you managae to get those numbers just curious

OpenStudy (qqstory):

yes.

OpenStudy (anonymous):

You either pick up a blue or orange in the first pick, then you pick a blue in the second pick. That is what I wrote up.

OpenStudy (anonymous):

and lol 35 + 54 = 89 and 89/87 is not equal to 5/8 :s

OpenStudy (qqstory):

i don't have the answer lol.

OpenStudy (anonymous):

the answer is 5/8

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@hamza_b23 , It seems that you understand my solution since yours is exactly the same as mine.

OpenStudy (anonymous):

i actually dont

OpenStudy (anonymous):

Read on A= Orange than Blue B= Blue than Blue We need \[ P[A \cup B] = P[ A] + P[B] = \frac{3\ 5}{8\ 7}+\frac{5\ 4}{8\ 7}= \frac 5 8 \]

OpenStudy (anonymous):

i am really sorry to break it down to you BUT (35 +54)/87 = 89/87 WHICH is greater than 1 HENCE not a probability and there is no way on earth that is EQUAL to 5/8 lol

OpenStudy (anonymous):

\[P[A \cup B] = P[ A] + P[B] = \frac 3 8 \frac 5 7+ \frac 5 8 \frac 4 7= \frac {15}{56}+ \frac {20}{56}=\frac {35}{56} = \frac 5 8 \]

OpenStudy (anonymous):

LOL O they are multiplied yes now it makes sense lol

OpenStudy (anonymous):

Can you explain how you got 5/8 as an answer?

OpenStudy (qqstory):

elisasaab ur smart.